OpenStudy (anonymous):

find the distance between points (1,0) and (3,2). leave your answer in simple radical form.

OpenStudy (anonymous):

\[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (radar):

\[\sqrt{(3-1)^{2}+(2-0)^{2}}\] Does that help any?

OpenStudy (radar):

\[\sqrt{2^{2}+2^{2}}\] for further assistance.

OpenStudy (radar):

\[\sqrt{8}=2\sqrt{2}\]

OpenStudy (radar):

That would be my final answer.

OpenStudy (anonymous):

ok thank you